OpenStudy (anonymous):

find the exact trigonometric ratios for the angle whose radian measures is given 11π/4

OpenStudy (anonymous):

first note that \[\frac{11\pi}{4}\] is coterminal with \[\frac{11\pi}{4}-2\pi=\frac{11\pi}{4}-\frac{8\pi}{4}=\frac{3\pi}{4}\] then locate \[\frac{3\pi}{4}\] on the unit circle on the attached cheat sheet to find all the trig functions at that piont

OpenStudy (anonymous):

thank you very much this is very helpful :)

OpenStudy (anonymous):

yw it is a very useful cheat sheet i think